If absorbance is 0.486 and the extinction coefficient, molar absorbtivity is 36000 per molar per cm at 414 nm. And you take 100 microlitre of the given solution in a 1 ml cuvette with 900 microlitre buffer. What is the concentration of the given stock of ABTS?
1
Expert's answer
2019-07-24T03:49:00-0400

Solution.

To find the concentration of the ABTS stock, it is necessary to write the Bouguer – Lambert – Beer Law and derive the concentration from this equation:


"A = \epsilon \times l \times C"

"C = \frac{A}{\epsilon \times l}"


"C = \frac{0.486}{36000 \times 1}"

"C = 1.35*10^{-5} M"

Answer:

"C = 1.35*10^{-5} M"